### v1

# Weakly supervised image segmentation for defect-based grading of fresh produce

Abstract
Implementing image-based machine learning in agriculture is often limited by scarce data and annotations, making it hard to achieve high-quality model predictions. This study tackles the issue of postharvest quality assessment of bananas in decentralized supply chains. We propose a method to detect and segment surface defects in banana images using panoptic segmentation to quantify defect size and number. Instead of time-consuming pixel-level annotations, we use weak supervision with coarse labels. A dataset of 476 smartphone images of bananas was collected under real-world field conditions and annotated for bruises and scars. Using the Segment Anything Model (SAM), a recently published foundation model for image segmentation, we generated dense annotations from coarse bounding boxes to train a segmentation model, significantly reducing manual effort while achieving a panoptic quality score of 77.6%. This demonstrates SAM’s potential for low-effort, accurate segmentation in agricultural settings with limited data.
keywords:
Machine Learning , Computer Vision , Food Quality , Postharvest
1 Introduction
Visual inspection of agricultural products plays a crucial role in postharvest supply chains. Traditionally, this process has relied heavily on manual inspection, which is labor-intensive, time-consuming, and prone to human error and inconsistency. The inspectors also need dedicated training, for example, to identify different defects, their size, and their severity. Typical defects for banana fruit are, for example, surface bruises and scars, insect damage, sunburn, chilling injuries, malformed fruits, or maturity stains. As a result, there has been a growing interest in leveraging advanced technologies such as computer vision and machine learning to enhance the accuracy and efficiency of visual inspection and grading of agricultural products (Yuvaraj et al., 2023; Barthwal et al., 2024). That way, inspectors with limited training could still accurately identify different defects and the severity of these defects in the field via a smartphone, which is very scalable, also to remote areas.
Machine learning-based computer vision models can establish an automated assessment procedure that increases the speed and consistency of grading. Traditionally, these AI models, however, require a considerably large and well-annotated training dataset to work well in real-life scenarios. However, large parts of the local, regional, and export postharvest supply chain are decentralized, non-automated, and not digitalized from farm to retail. Therefore, stakeholders often do not have large annotated datasets readily available for their specific use case, defects, and cultivars that could be used to set up these AI models (De-Arteaga et al., 2018). Recent computer vision research shows a trend toward foundation models that are pre-trained on large (often unlabeled) datasets, enabled by advances in self-supervised learning (Balestriero et al., 2023). These models can be more easily adapted to complex downstream tasks by requiring less fine-tuning data compared to the previous generation of models and, therefore, facilitate the implementation of machine learning solutions in low data regimes (Knott et al., 2023). Most foundation models can be adapted to various tasks in a specific data modality (e.g., text, image, or time series) or across modalities (e.g., Li et al., 2024; Zhang et al., 2023). Other foundation models are designed to excel in a certain task or domain problem within a modality (e.g., Israel et al., 2023; Shah et al., 2023; Lu et al., 2024).
In this work, we aim to demonstrate that visual foundational models can be used to detect defects based on a relatively low number of images and with coarse annotations only. We address a common type of visual inspection: recognizing surface defects such as blemishes caused by external impacts (Abbas et al., 2019). Our study focuses on banana fruits and images taken via non-static cameras (e.g., smartphones) in the field and in sorting and packing facilities. We identified two key indicators relevant for grading: the number and size of blemishes visible on a bunch. In addition, the type of the defect is of secondary interest as it can help supply chain operators identify the cause of the defect. Such a use case is typically approached via Image Segmentation, a subcategory of Computer Vision that requires training data with dense image annotations, meaning each pixel has to be associated with a category and/or instance, traditionally leading to a tedious annotation process. To address this, our method utilizes the Segment Anything Model (SAM, Kirillov et al., 2023), a visual foundation model that specializes in image segmentation. We use SAM to generate dense image annotations from coarse high-level annotations (bounding boxes) and use them to train another image segmentation model that identifies, classifies, and segments banana surface defects in an image. We demonstrate that this method works well on a small dataset of 476 coarsely annotated images.
2 Materials and Methods
2.1 Data collection
We collected 476 images of single banana bunches containing at least one visible bruise or scar in three different sorting and packing facilities in India located in Tembhurni (Maharashtra state), Raver (Maharashtra state), and Kamrej (Gujarat state). The images were taken with different smartphone cameras between harvesting and packing. Due to the small dataset size, these images are split into five subsets used for cross-validation. This means that we evaluate all results on five different train-validation splits to ensure the robustness of our results. To ensure independent data samples, each bunch was only photographed once. The photos have realistic, non-standardized properties: the proximity to the fruits, lighting conditions, image sharpness, and resolution can vary. Also, there can be other bananas and objects occasionally in the background. Such conditions are typical for images taken in supply chains by inspectors, so we wanted our method to accommodate that. We aim to first automatically detect, segment, and classify defects using a machine learning model. In the second step, we derive the number of visible surface defects and their relative size from these predictions.
There are different principal approaches to defining target labels in image segmentation. One can assign each pixel to a general class (“semantic segmentation”) or to both a class and an instance (“instance segmentation”). The choice depends on the use case. This work uses a mixed approach called “panoptic segmentation”, which we will justify in the following paragraphs. See Section 2.2 for a more detailed explanation of the segmentation concepts.
Figure 1: Defect examples. Example images of banana surface defects we aim to detect in this study. Bruises usually result from a dull impact, while scars are caused due to a sharp impact. Old defects are usually darker than new defects as they oxidize over time.
Defect annotations
The images were then annotated by a single expert using the open-source tool “LabelStudio” (Tkachenko et al., 2020). As we are interested in the number of defects, we annotated defects on an instance level, meaning that we annotated one bounding box per defect. While a key property of our approach is to only use bounding boxes and generate instance masks via SAM, we also annotate defect masks for validation purposes. Specifically, each defect was one of four categories: “Old Bruise”, “New Bruise”, “Old Scar”, and “New Scar”.
Bruises typically result from blunt force or pressure, which can arise from various factors such as handling, packing, impacts, or weight. Scars, on the other hand, are caused by sharp contact due to reasons like friction from leaves or stems, cuts from knives, crates, rough handling, laborers’ nails, or the sharp tips and crowns of neighboring bunches. Old defects are usually darker than new defects as they oxidize over time (see Figure 1). While the automated detection of generic defects without classification is already an impactful use case, we chose to add this categorization as it helps supply chain managers trace back the cause of the damage in the supply chain. It is important to note that there are cases that cannot be unambiguously assigned to one of those categories, even by expert annotators. However, a sub-goal of this work is to assess to which degree an automated classification is possible. It is worth mentioning that there are plenty of other relevant surface defects, reducing the quality of banana crops that are not included in this study, e.g., insect damage, speckles, sunburn, maturity stains, malformed fruits, broken necks, or chemical residues.
Banana annotations
To obtain relative defect sizes, we also annotated segmentation masks of banana fruits. We decided to only annotate bananas semantically on a pixel level, omitting instance information for two reasons: First, as images often contain multiple bunches that are partly cropped or occluded, it is challenging to differentiate instances. Second, our use case does not require us to identify instances, as we are not interested in counting single bananas or bunches. We would rather distinguish focused foreground bunches from those that may appear in the background and, therefore, annotate two types of pixel classes: “Foreground Banana” and “Background Banana”. All pixels not annotated as a banana of defect are assigned the “Background” class. We annotated banana masks using LabelStudio and Segment Anything point prompts (see Section 2.3 for more details).
Figure 2: Overview of our approach. We utilize the Segment Anything Model (SAM), a promptable visual foundation model for image segmentation, to generate dense annotations (pixel-wise class and instance labels) from coarse annotations (bounding boxes and reference points) without any model training involved. These dense annotations otherwise require tedious hand-annotation. Using these newly generated labels, we train a panoptic segmentation model to identify surface defects on banana fruits, specifically bruises and scars. Additionally, the model can differentiate between foreground banana fruits and those in the background. This enables us to determine the number and size of visible surface defects from photographs of banana fruits. We validate our approach using a dataset of 476 images.
2.2 Image segmentation
Image segmentation uses computer vision techniques to analyze and understand images on a pixel level. In this section, we describe different paradigms of image segmentation and their corresponding evaluation metrics that we use to evaluate the quality of our models.
2.2.1 Semantic Segmentation
“Semantic segmentation” is concerned with classifying each pixel of an image into a predefined category (e.g., background, banana fruit, defect) without distinguishing between individual objects (instances) of the same class. (Ronneberger et al., 2015; Chen et al., 2017; Xie et al., 2021)
Semantic segmentation maps are evaluated against a groundtruth using the Intersection-over-Union (IoU) score for a single class or the mean IoU (mIoU) score for all $N$ classes in a dataset (Everingham et al., 2012; Lin et al., 2014). Eqs. 1 and 2 show the definition of those metrics where $G$ refers to the set of groundtruth pixels that are assigned to a class and $P$ to the set of prediction pixels accordingly.
$IoU(G,P)$ $=\frac{|G\cap P|}{|G\cup P|}$ (1)
$mIoU$ $=\frac{1}{N}\sum_{i=1}^{N}IoU_{i}$ (2)
2.2.2 Instance Segmentation
Instance segmentation goes a step further by categorizing each pixel and identifying each object instance of the same class separately. This is realized by assigning each pixel to one of potentially multiple objects and to a semantic class simultaneously. This technique is essential for tasks requiring counting or differentiating between items. It is considered more complex, combining object detection with pixel classification. (He et al., 2017; Lin et al., 2017)
In terms of evaluation, instance segmentation is equivalent to object detection, which in machine learning describes the localization of instances via rectangular bounding boxes. The only difference is that the overlap of instances is defined by masks rather than boxes. There are several widely recognized metrics that compare a set of groundtruth bounding boxes or instance masks to a set of predicted bounding boxes or instance masks. Generally, a predicted and an annotated instance are matched if their localization exceeds a certain IoU threshold. As a first step, we calculate the precision $P$ and recall $R$ as follows:
$P$ $=\dfrac{TP}{TP+FP}$ (3)
$R$ $=\dfrac{TP}{TP+FN}$ (4)
where $TP$ is the number of true positives, $FP$ is the number of false positives, and $FN$ is the number of false negatives. By definition, precision and recall are inversely related but must concur effectively to ensure that both relevant and correct instances are identified. The primary evaluation metric for object detection is Average Precision ($AP$) for a single class and mean Average Precision ($mAP$) for multiple classes. We follow established practices and report three different variations:
  • $AP@50$: $AP$ at a $IoU$ threshold of $0.50$ as the most common metric in many benchmarks (e.g., Everingham et al., 2012),
  • $AP@75$: $AP$ at a $IoU$ threshold of $0.75$ as a more strict metric,
  • $AP$ (or $AP^{IoU=.50:.05:.95}$): $AP$ averaged over ten equidistant $IoU$ thresholds between $0.50$ and $0.95$ as proposed by Lin et al. (2014).
The average recall is another evaluation metric used to measure the assertiveness of object detectors for a given class (Padilla et al., 2021). Similar to $AP$, the $AR$ metric evaluates a range of IOU thresholds. We use the equation provided by Hosang et al. (2015) to calculate $AR$ from a discrete set of thresholds based on the total number of ground truth instances $G$ in an image:
$AR=\dfrac{2}{G}\sum_{i=1}^{G}max(IoU_{i}-0.5,0)$ (5)
2.2.3 Panoptic Segmentation
In scenarios where one requires both instance-level information for some classes that are countable (widely referred to as “things”) and semantic information about other non-countable components (referred to as “stuff”), these two concepts need to be combined into an approach called “Panoptic Segmentation”. This approach seeks to overcome the limitations of its predecessors by delivering a comprehensive segmentation map that includes both class labels and instance identification. (Cheng et al., 2021, 2022; Kirillov et al., 2019a, b)
In our study, we declare defects as “things” and bananas as “stuff” as we are not interested in distinguishing single banana fruits. Thus, we simplify the machine learning task.
Kirillov et al. (2019b) propose Panoptic Quality ($PQ$) as a comprehensive metric for panoptic image segmentation (Eq. 6). The intuition behind $PQ$ is to provide a unified measure that captures the accuracy of detecting and correctly segmenting individual objects (defects in our case) as well as the overall segmentation quality of the entire scene. $PQ$ combines precision (how many predicted segments are correct) and recall (how many true segments are detected) into a single score, rewarding models that can accurately identify object boundaries and classify them correctly.
$PQ=\frac{\sum_{(p,g)\in TP}\text{IoU}(p,g)}{|TP|+\frac{1}{2}|FP|+\frac{1}{2}|FN|}$ (6)
In this work, we use Panoptic Quality as the primary evaluation metric to compare how well trained image segmentation models perform.
2.3 Segment Anything Model (SAM)
The Segment Anything Model (SAM) is a recently published foundation model for image segmentation (Kirillov et al., 2023). It was trained on a large-scale dataset comprising diverse images from various domains, including natural scenes, urban environments, medical imagery, and other contexts. It is the first model of its kind specifically tailored for generalized image segmentation and promises to facilitate the implementation of domain-specific segmentation models.
Figure 2 (left) shows a high-level overview of SAM’s components. The model consists of three major components: (1) a vision transformer (Dosovitskiy et al., 2020) pre-trained with the Masked Autoencoder method (He et al., 2021) that serves as an image encoder; (2) a prompt encoder that allows the user to specify bounding boxes or points to localize the region of interest in an image [^1] In their paper, the authors of SAM mention open vocabulary text as input prompts. However, they did not publish their code for this feature as the results are not as good as for other prompts (REF needed). Therefore, we do not consider this option in this research.; (3) a mask decoder that generates segmentation masks based on the concatenated image and prompt embeddings. This approach yields impressive results for zero-shot instance segmentation in various domains. However, compared to conventional segmentation models, it comes with two disadvantages that require further adaptations or post-processing to enable SAM for a typical image segmentation use case: Firstly, SAM requires weak labels such as bounding boxes or point coordinates as input prompts. Secondly, the output masks are uncategorized, which hinders zero-shot usage of SAM in a multi-class setting.
SAM2 (Ravi et al., 2024) is a recent extension of SAM to the video domain. Although videos are not the focus of this work, SAM2 has demonstrated that their video-based pre-training enhances zero-shot mask generation for images as well. In our study, we assess both SAM and SAM2 using a ViT-L backbone for zero-shot mask generation of banana fruits and defects.
2.4 Experimental setup
Figure 2 shows an overview of our approach. The process starts with an image dataset with coarsely annotated labels: bounding boxes to locate and assign labels to defects and sets of points that identify foreground and background bananas. We utilize a promptable foundation model for image segmentation (SAM) to generate fine-grained pixel-wise annotations of each image in the dataset based on the coarse hand annotations. It is important to note that SAM is not trained in this procedure but is used here in a zero-shot manner to infer dense annotations only. These newly created annotations are used to train a panoptic segmentation model in a standard supervised machine learning workflow. The panoptic segmentation model we use is a Maskformer (Cheng et al., 2021) model pre-trained on the ADE20k dataset (Zhou et al., 2017). Further details on model training, such as hyperparameters and data augmentation, can be found in A.
Cross-validation
Given the relatively small size of our training dataset, it is important to ensure that our reported results are robust. To achieve this, we use five-fold cross-validation instead of a static validation split. This approach involves dividing the dataset into five equally sized parts. We repeat each experiment five times, with one partition serving as the validation set while the remaining four partitions are used to train the model in each iteration. The evaluation metrics are then reported as the mean and standard deviation across these five experiments.
Validating SAM masks.
Qualitative analysis indicated that SAM-generated defect masks are generally plausible. However, to quantify the impact of training on SAM-generated masks compared to using hand-annotated pixel-wise masks, we repeat our experiments with different combinations of mask sources for training and validation data: (a) both training and validation use hand-annotated defect masks; (b) both training and validation use SAM-generated defect masks; and (c) training uses SAM-generated masks while validation uses hand-annotated masks. By comparing the evaluation metrics across these three variations, we can quantify the trade-off between the time-saving benefits of automated mask generation and the quality of the resulting output.
2.5 Postprocessing
In some cases, it can be ambiguous whether to treat closely located defects as a single defect or multiple defects. To address this and ensure more consistent defect separation, we introduce a simple algorithm-based postprocessing method with the following three steps: (1) Combine all defect instances of the same type into a single binary map. (2) Use a connected components algorithm (OpenCV, 2015) to generate new instances where no pixels are connected. 3) Apply a dilation operation to expand all instance masks by $d$ pixels (we choose $d=5$) in all directions. This expanded auxiliary mask is then used to identify overlapping instances, which are subsequently merged. In other words, defect instances of the same class within a distance of $2d$ pixels are merged.
Figure 3: Distribution of IoU values when comparing annotated and SAM-generated defect masks. We generally observe a high overlap, while SAM2 masks are closer aligned to the hand-annotated ones.
2.6 Deriving number and size of defects
In the context of quality estimation, we are interested in two practical indicators that can be derived from an image’s panoptic prediction: The number and size of defects. While the first can be estimated by counting the number of defect instances, an accurate metrical size estimation in a two-dimensional image is impossible due to unknown scale and distortions. Thus, we use the relative defect size, simplified as the number of visible defect pixels divided by the number of visible foreground banana pixels, as a proxy.
3 Results and Discussion
3.1 Comparison of annotated and generated masks
Our approach uses SAM to generate dense segmentation masks in a zero-shot setting (i.e., without pre-training). To assess the quality of these masks, we first compare the masks produced by different SAM models (SAM and SAM2) with hand-annotated ground truth masks (Figure 3). Overall, we observe a high overlap between the hand-annotated defect masks and those generated by both SAM models. Notably, SAM2 exhibits even greater accuracy than SAM, with 90.5% of its masks achieving an IoU score of 0.5 or higher. Based on these findings, we conclude—and later demonstrate—that our zero-shot mask generation approach is both valid and capable of producing segmentation results comparable to hand-annotated masks.
Figure 4: Hand-annotated vs. estimated defect counts per image. We predict the exact number of defects 36.5% of the time (green squares). 78.3% of the time, we predict correctly within a ±$\pm 1$ tolerance (orange squares). Generally, our model tends to predict more defects in the images compared to the annotations.
Figure 5: Annotated vs estimated defect sizes. We pairwise match those defect instances with the highest IoU agreement between annotation and estimation (minimum 0.5) and calculate their sizes relative to the corresponding foreground banana masks. Relative size prediction is generally accurate with a Pearson correlation of $r=0.96,n=829$. Each blue dot is one pair of defects, the orange line is a fitted regression line, black line is $x=y$ diagonal.
3.2 Evaluation with image segmentation metrics
Next, we assess image segmentation results on validation images using the metrics outlined in Section 2.2. Table 1 presents these metrics across various experimental setups: In the first experiment, we used hand-annotated defect masks for both training and validation, achieving an average panoptic quality (PQ) of 77.5% after applying postprocessing. This serves as our baseline for assessing how segmentation quality changes when replacing hand-annotated masks with SAM-generated ones. In the second experiment, we employ SAM-generated masks for both training and validation. The results show comparable PQ scores of 77.3% (SAM) and 77.6% (SAM2). In the final experiment, SAM-generated masks are used for training, while hand-annotated masks are used for validation. Here, the PQ scores are approximately $\mathord{\sim}0.5\%$ lower for both SAM and SAM2 compared to when hand-annotated masks are used for training. In conclusion, while SAM-generated masks are not fully equivalent to hand-annotated ones—occasionally failing to capture certain edge cases—they still produce metrics that are sufficiently close to be considered a viable low-effort alternative: For our dataset, we estimated an annotation time of approx. 15 seconds per image for coarse annotations and 3-5 minutes per image for pixel-wise annotations. An additional interpretation of these results is that there is a slight but consistent difference in how our human annotator and SAM delineate defect boundaries. Nonetheless, the Maskformer model demonstrates the ability to learn from both annotation styles effectively. Lastly, it is evident that our postprocessing method is effective for improving defect masks.
3.3 Qualitative evaluation
The results can be qualitatively analyzed through the example visualizations in Figure 6 showing the segmentation using a single generic defect class. In general, the masks appear sharp, particularly for bananas, and the model reliably distinguishes between fruits in the foreground or center of the image and those in the background. While the detected defects do not always align with the annotations, they are generally plausible. The model often identifies blemishes that were not annotated, likely because they were deemed too minor to be significant. This underscores the challenge of defining ground-truth defect annotations in these images.
3.4 Number and size of defects
Panoptic segmentation masks can simply be used to derive the number and size of defects. We compare the number of defects detected per image (Figure 4) and the relative defect sizes (Figure 5) by our best model (x-axes) and the human annotations (y-axes).
The model accurately matches the exact number of defects in 36.5% of cases and is off by one in 78.3% of cases. This further highlights the issue of ambiguous ground truth, as discussed earlier.
For relative size analysis, we pairwise match those defect instances with the highest IoU agreement between annotation and estimation. If no agreement with at least 0.5 IoU can be established, we consider predicted defects non-matchable and exclude them from this analysis. Relative sizes are calculated by dividing the number of pixels of a specific defect instance by the aggregated number of pixels assigned to either the foreground banana class or any defect. We generally see a high agreement in predicted vs. annotated relative sizes ($r=0.96)$, while the model tends to slightly underestimate sizes, visible by the different slopes of the regression line (orange) and $x=y$ (black dashed) lines in Figure 5.
Table 1: Main results with a single defect category. All results are reported via their mean and standard deviation over the five different cross-validation splits. We compare different setups where target defect masks are either annotated (“Anno.”) or generated by SAM/SAM2. Expectedly, when Maskformer is trained on SAM-generated masks, the defect IoU is slightly lower when evaluating annotated masks. However, recall and precision are comparable to the fully annotated setting. We also demonstrate that our postprocessing (“PP”) significantly improves precision and recall of defects. Banana segmentation quality is slightly better in the fully annotated setting ($\mathord{\sim}0.81$ vs $\mathord{\sim}0.80$ mIoU). The configuration highlighted in blue is used for subsequent visualizations.
Defect masks Defects Fruits Overall Model PP train val AP AP@50 AP@75 AR IoU IoU FG IoU BG mIoU PQ Maskformer Anno. Anno. .$133\pm.022$$243\pm.027$$134\pm.027$$256\pm.036$$593\pm.032$$952\pm.005$$732\pm.034$$812\pm.010$$758\pm.010$ Maskformer Anno. Anno. .$220\pm.019$$422\pm.014$$215\pm.033$$412\pm.041$$593\pm.032$$952\pm.005$$732\pm.034$$813\pm.010$$775\pm.007$ Maskformer SAM Anno. .$128\pm.030$$234\pm.043$$128\pm.040$$238\pm.025$$546\pm.040$$951\pm.006$$731\pm.052$$800\pm.011$$752\pm.011$ Maskformer SAM Anno. .$205\pm.017$$392\pm.040$$197\pm.013$$380\pm.022$$548\pm.039$$951\pm.006$$731\pm.052$$801\pm.011$$770\pm.005$ Maskformer SAM SAM .$137\pm.030$$241\pm.045$$4.1$034 .$250\pm.027$$561\pm.038$$951\pm.006$$731\pm.052$$804\pm.012$$755\pm.010$ Maskformer SAM SAM .$220\pm.016$$404\pm.044$$2$016 .$399\pm.023$$562\pm.037$$951\pm.006$$731\pm.052$$804\pm.011$$773\pm.005$ Maskformer SAM2 Anno. .$121\pm.028$$227\pm.047$$121\pm.036$$238\pm.034$$5.5$025 .$950\pm.005$$720\pm.041$$799\pm.015$$753\pm.011$ Maskformer SAM2 Anno. .$201\pm.022$$395\pm.018$$184\pm.030$$392\pm.020$$554\pm.025$$950\pm.005$$720\pm.041$$799\pm.015$$770\pm.013$ Maskformer SAM2 SAM2 .$134\pm.039$$229\pm.050$$139\pm.054$$249\pm.044$$579\pm.028$$950\pm.005$$720\pm.041$$805\pm.015$$757\pm.012$ Maskformer SAM2 SAM2 .$225\pm.035$$405\pm.029$$221\pm.056$$415\pm.034$$579\pm.028$$950\pm.005$$720\pm.041$$805\pm.015$$776\pm.012$
3.5 Defect Categorization
We also conducted experiments using four distinct defect types instead of a single defect class (see B for details). In this setup, the model must not only detect and segment defects but also categorize them into one of four pre-defined classes. Our results clearly show that while the detection and segmentation of bruises and scars work well, a reliable categorization into one of the four predefined classes is not possible with the current approach and dataset. We hypothesize that this limitation stems from one or more of the following factors:
  1. 1.
    Ambiguous annotation: The classification of defects across the four predefined categories may be subjective to a large degree. The distinction between “old” and ”new“ defects is not always clear-cut, as the transition is gradual. Additionally, distinguishing bruises from scars can be ambiguous, especially for non-experts (see Figure 1).
  2. 2.
    Image resolution: We use $1024^{2}$ pixel resolution for images. While this is generally considered high for machine learning tasks, defects can be small and, thus, only be represented by a few pixels, making them harder to categorize.
  3. 3.
    The four defect types are unevenly represented in our dataset (37/82/387/834). A more balanced distribution of categories and a larger number of defect samples are likely to improve categorization accuracy. We recommend collecting more examples from the two underrepresented classes for future work.
Figure 6: Example visualizations of annotated vs predicted masks using a single defect category. Left: Input Image, Mid: Annotation (SAM), Right: Maskformer Prediction. Segments are color-coded as follows: Foreground Banana, Background Banana, Defect,
4 Conclusion and Outlook
In this work, our objective was to have an efficient way to implement machine learning to count the number of defects on banana fruit and have their relative sizes. We demonstrated the use of a panoptic image segmentation model to tackle the detection and pixel-wise segmentation of bananas and surface blemishes in images. Our approach significantly reduces the effort required for hand annotation by relying on coarse labels. Specifically, we utilize a pre-trained foundation model for image segmentation (the Segment Anything Model) to generate the dense annotations necessary for training the panoptic model. This method reduces the manual annotation workload by at least a factor of 10 while maintaining high-quality results: the measured Panoptic Quality drops from 77.5% to 77.0%) w.r.t. the annotated targets.
Our results generally demonstrate reliable and plausible segmentation of both foreground and background bananas, as well as a single generic defect category. However, when compared to human annotations, the model matches the exact number of defects in only 36.5% of cases and is off by up to one defect in 78.3% of cases. While some of this discrepancy can be attributed to model performance, it is more likely due to the target annotations. Unlike for other machine learning applications where target labels are often close to a “ground truth”, the identification of significant blemishes is to a certain degree subjective. An improvement on the labeling side could be to employ multiple annotators and measure their level of agreement. The benefits of using (noisy) annotations from multiple annotators for samples that are non-trivial to label have been explored in the machine learning literature (Welinder et al., 2010; Sheng et al., 2008). Thus, we recommend that future work carries out a more sophisticated annotation process with multiple expert annotators. The disagreement between annotators could then serve as a baseline for evaluating the performance of future machine learning models.
In addition, we demonstrated that with our current method and dataset, it is not possible to reliably categorize defects into one of the four predefined classes, namely new scars, old scars, new bruises, and old bruises. A variable ground truth is likely a contributing factor, along with the skewed class distribution and the potentially low image resolution for small defect sizes. Addressing these challenges will require further investigation in future work to improve the reliability of defect categorization.
Our method demonstrates the potential for dense image segmentation in low-data scenarios and can be broadly applied to any commodity with countable, localizable defects. However, for defects that are non-countable—such as speckles, sunburn, or chemical residues—a different machine learning paradigm (such as classification) would be more applicable.
Acknowledgments
This work was partially funded by the project “Scaling up Your Virtual Cold Chain Assistant”, commissioned by the German Federal Ministry of Economic Cooperation and Development (BMZ) on behalf of the German International Cooperation Agency (GIZ).
Supplementary materials
The source code of all experiments, as well as pre-trained models, can be accessed on GitHub: https://github.com/manuelknott/banana-defect-segmentation.
Supplementary Materials
Appendix A Implementation details
All images were resized and padded to $1024^{2}$ pixels resolution. Training images were augmented with a 50% Random Horizontal Flip and with additional random color-based augmentations (see Table S.1) to increase model robustness.
Table S.2 shows the hyperparamters used for training the Maskformer models. We evaluated the model every five epochs on the validation set and saved the model with the highest Panoptic Quality. All trainings were done on a single NVIDIA RTX A5000 GPU with 24GB memory.
Table S.1: Random color augmentations applied to training images. Values are uniformly sampled from the specified ranges.
Property Sampling range
brightness [0.9, 1.1]
contrast [0.9, 1.1]
saturation [0.9, 1.1]
hue [-0.05, 0.05]
Table S.2: Hyperparameters for Maskformer training.
Batch size 2
Epochs 100
Evaluation frequency every 5 epochs
Optimizer Adam (Kingma, 2014)
Learning rate $5\times 10^{-5}$
Learning rate schedule constant
Appendix B Results using four defect categories
Table S.3: Results using multiple defect categories. It is evident that our models are unable to categorize defect types. Most likely due to ambiguous annotations and/or limited training data. The configuration highlighted in blue is used for the visualizations in Figure S.1.
Defects Defect masks Old Bruise New Bruise Old Scar New Scar Overall Model PP train val AP IoU AP IoU AP IoU AP IoU mAP mIoU PQ Maskformer Anno. Anno. .$027\pm.060$$011\pm.024$$058\pm.024$$186\pm.107$$039\pm.035$$229\pm.048$$063\pm.026$$376\pm.077$$047\pm.020$$493\pm.018$$471\pm.013$ Maskformer Anno. Anno. .$009\pm.020$$051\pm.061$$063\pm.013$$115\pm.067$$030\pm.021$$199\pm.062$$034\pm.013$$123\pm.061$$034\pm.009$$447\pm.028$$442\pm.021$ Maskformer SAM Anno. .$0.000\pm.0$000 .$0.000\pm.0$000 .$058\pm.031$$166\pm.078$$031\pm.019$$243\pm.038$$064\pm.035$$351\pm.067$$038\pm.014$$488\pm.012$$462\pm.008$ Maskformer SAM Anno. .$0.000\pm.0$000 .$003\pm.006$$071\pm.024$$079\pm.046$$024\pm.018$$178\pm.055$$4.0$014 .$122\pm.035$$034\pm.010$$433\pm.014$$440\pm.007$ Maskformer SAM SAM .$0.000\pm.0$000 .$0.000\pm.0$000 .$061\pm.034$$173\pm.081$$035\pm.020$$247\pm.042$$069\pm.036$$359\pm.067$$041\pm.014$$490\pm.013$$467\pm.007$ Maskformer SAM SAM .$0.000\pm.0$000 .000±.001plus-or-minus.000.001.000$\pm$.$0$001 .$077\pm.026$$083\pm.049$$029\pm.024$$180\pm.058$$043\pm.014$$125\pm.036$$037\pm.010$$434\pm.016$$442\pm.006$ Maskformer SAM2 Anno. .$0.000\pm.0$000 .$0.000\pm.0$000 .$047\pm.019$$144\pm.082$$039\pm.038$$235\pm.086$$062\pm.014$$394\pm.052$$037\pm.007$$490\pm.021$$467\pm.009$ Maskformer SAM2 Anno. .$0.000\pm.0$000 .$021\pm.031$$064\pm.033$$127\pm.083$$027\pm.019$$168\pm.064$$033\pm.012$$094\pm.038$$031\pm.007$$437\pm.017$$437\pm.017$ Maskformer SAM2 SAM2 .$0.000\pm.0$000 .$0.000\pm.0$000 .$049\pm.019$$152\pm.086$$055\pm.065$$242\pm.090$$067\pm.012$$405\pm.053$$043\pm.014$$494\pm.022$$472\pm.011$ Maskformer SAM2 SAM2 .$0.000\pm.0$000 .$020\pm.030$$067\pm.036$$133\pm.086$$038\pm.034$$170\pm.066$$036\pm.013$$098\pm.039$$035\pm.003$$439\pm.018$$442\pm.019$
Figure S.1: Example visualizations of annotated vs. predicted masks using four defect categories. Left: Input Image, Mid: Annotation, Right: Maskformer Predictions (SAM2). Segments are color-coded as follows: Foreground Banana, Background Banana, Old Bruise, Old Scar, New Bruise, New Scar.

### v2

@@
-Weakly supervised image segmentation for defect-based grading of fresh produce
+Weakly Supervised Panoptic Segmentation for Defect-Based Grading of Fresh Produce
 Abstract
-Implementing image-based machine learning in agriculture is often limited by scarce data and annotations, making it hard to achieve high-quality model predictions.
-This study tackles the issue of postharvest quality assessment of bananas in decentralized supply chains.
-We propose a method to detect and segment surface defects in banana images using panoptic segmentation to quantify defect size and number.
-Instead of time-consuming pixel-level annotations, we use weak supervision with coarse labels.
-A dataset of 476 smartphone images of bananas was collected under real-world field conditions and annotated for bruises and scars.
-Using the Segment Anything Model (SAM), a recently published foundation model for image segmentation, we generated dense annotations from coarse bounding boxes to train a segmentation model, significantly reducing manual effort while achieving a panoptic quality score of 77.6%.
-This demonstrates SAM’s potential for low-effort, accurate segmentation in agricultural settings with limited data.
-keywords:
-Machine Learning , Computer Vision , Food Quality , Postharvest
+Visual inspection for defect grading in agricultural supply chains is crucial but traditionally labor-intensive and error-prone.
+Automated computer vision methods typically require extensively annotated datasets, which are often unavailable in decentralized supply chains.
+We address this challenge by evaluating the Segment Anything Model (SAM) to generate dense panoptic segmentation masks from sparse annotations.
+These dense predictions are then used to train a supervised panoptic segmentation model.
+Focusing on banana surface defects (bruises and scars), we validate our approach using 476 field images annotated with 1440 defects.
+While SAM-generated masks generally align with human annotations, substantially reducing annotation effort, we explicitly identify failure cases associated with specific defect sizes and shapes.
+Despite these limitations, our approach offers practical estimates of defect number and relative size from panoptic masks, underscoring the potential and current boundaries of foundation models for defect quantification in low-data agricultural scenarios.
+GitHub: https://github.com/manuelknott/banana-defect-segmentation
+footnotetext: Accepted as a paper to the 6th International Workshop on Agriculture-Vision: Challenges & Opportunities for Computer Vision in Agriculture in conjunction with IEEE/CVF CVPR 2025.
 1 Introduction
 Visual inspection of agricultural products plays a crucial role in postharvest supply chains.
-Traditionally, this process has relied heavily on manual inspection, which is labor-intensive, time-consuming, and prone to human error and inconsistency.
-The inspectors also need dedicated training, for example, to identify different defects, their size, and their severity.
-Typical defects for banana fruit are, for example, surface bruises and scars, insect damage, sunburn, chilling injuries, malformed fruits, or maturity stains.
-As a result, there has been a growing interest in leveraging advanced technologies such as computer vision and machine learning to enhance the accuracy and efficiency of visual inspection and grading of agricultural products (Yuvaraj et al., 2023; Barthwal et al., 2024).
-That way, inspectors with limited training could still accurately identify different defects and the severity of these defects in the field via a smartphone, which is very scalable, also to remote areas.
-Machine learning-based computer vision models can establish an automated assessment procedure that increases the speed and consistency of grading.
-Traditionally, these AI models, however, require a considerably large and well-annotated training dataset to work well in real-life scenarios.
+Particularly, fresh produce is often graded based on visible blemishes and defects.
+Traditionally, this process relies on manual inspection—an approach that is labor-intensive, time-consuming, and prone to inconsistency and human error.
+Additionally, inspectors require specialized training to recognize various types of defects, assess their size, and determine their severity.
+For those reasons, machine learning-based computer vision models have become increasingly popular for automated assessment, increasing the speed and consistency of grading [28, 3].
+These deep learning models traditionally require a considerably large and well-annotated training dataset to work well and robustly in real-world scenarios.
 However, large parts of the local, regional, and export postharvest supply chain are decentralized, non-automated, and not digitalized from farm to retail.
-Therefore, stakeholders often do not have large annotated datasets readily available for their specific use case, defects, and cultivars that could be used to set up these AI models (De-Arteaga et al., 2018).
-Recent computer vision research shows a trend toward foundation models that are pre-trained on large (often unlabeled) datasets, enabled by advances in self-supervised learning (Balestriero et al., 2023).
-These models can be more easily adapted to complex downstream tasks by requiring less fine-tuning data compared to the previous generation of models and, therefore, facilitate the implementation of machine learning solutions in low data regimes (Knott et al., 2023).
-Most foundation models can be adapted to various tasks in a specific data modality (e.g., text, image, or time series) or across modalities (e.g., Li et al., 2024; Zhang et al., 2023).
-Other foundation models are designed to excel in a certain task or domain problem within a modality (e.g., Israel et al., 2023; Shah et al., 2023; Lu et al., 2024).
-In this work, we aim to demonstrate that visual foundational models can be used to detect defects based on a relatively low number of images and with coarse annotations only.
-We address a common type of visual inspection: recognizing surface defects such as blemishes caused by external impacts (Abbas et al., 2019).
-Our study focuses on banana fruits and images taken via non-static cameras (e.g., smartphones) in the field and in sorting and packing facilities.
-We identified two key indicators relevant for grading: the number and size of blemishes visible on a bunch.
-In addition, the type of the defect is of secondary interest as it can help supply chain operators identify the cause of the defect.
-Such a use case is typically approached via Image Segmentation, a subcategory of Computer Vision that requires training data with dense image annotations, meaning each pixel has to be associated with a category and/or instance, traditionally leading to a tedious annotation process.
-To address this, our method utilizes the Segment Anything Model (SAM, Kirillov et al., 2023), a visual foundation model that specializes in image segmentation.
-We use SAM to generate dense image annotations from coarse high-level annotations (bounding boxes) and use them to train another image segmentation model that identifies, classifies, and segments banana surface defects in an image.
-We demonstrate that this method works well on a small dataset of 476 coarsely annotated images.
-2 Materials and Methods
-2.1 Data collection
-We collected 476 images of single banana bunches containing at least one visible bruise or scar in three different sorting and packing facilities in India located in Tembhurni (Maharashtra state), Raver (Maharashtra state), and Kamrej (Gujarat state).
-The images were taken with different smartphone cameras between harvesting and packing.
-Due to the small dataset size, these images are split into five subsets used for cross-validation.
-This means that we evaluate all results on five different train-validation splits to ensure the robustness of our results.
-To ensure independent data samples, each bunch was only photographed once.
-The photos have realistic, non-standardized properties: the proximity to the fruits, lighting conditions, image sharpness, and resolution can vary.
-Also, there can be other bananas and objects occasionally in the background.
-Such conditions are typical for images taken in supply chains by inspectors, so we wanted our method to accommodate that.
-We aim to first automatically detect, segment, and classify defects using a machine learning model.
-In the second step, we derive the number of visible surface defects and their relative size from these predictions.
-There are different principal approaches to defining target labels in image segmentation.
-One can assign each pixel to a general class (“semantic segmentation”) or to both a class and an instance (“instance segmentation”).
-The choice depends on the use case.
-This work uses a mixed approach called “panoptic segmentation”, which we will justify in the following paragraphs.
-See Section 2.2 for a more detailed explanation of the segmentation concepts.
-Figure 1: Defect examples.
-Example images of banana surface defects we aim to detect in this study.
-Bruises usually result from a dull impact, while scars are caused due to a sharp impact.
-Old defects are usually darker than new defects as they oxidize over time.
-Defect annotations
-The images were then annotated by a single expert using the open-source tool “LabelStudio” (Tkachenko et al., 2020).
-As we are interested in the number of defects, we annotated defects on an instance level, meaning that we annotated one bounding box per defect.
-While a key property of our approach is to only use bounding boxes and generate instance masks via SAM, we also annotate defect masks for validation purposes.
-Specifically, each defect was one of four categories: “Old Bruise”, “New Bruise”, “Old Scar”, and “New Scar”.
-Bruises typically result from blunt force or pressure, which can arise from various factors such as handling, packing, impacts, or weight.
-Scars, on the other hand, are caused by sharp contact due to reasons like friction from leaves or stems, cuts from knives, crates, rough handling, laborers’ nails, or the sharp tips and crowns of neighboring bunches.
-Old defects are usually darker than new defects as they oxidize over time (see Figure 1).
-While the automated detection of generic defects without classification is already an impactful use case, we chose to add this categorization as it helps supply chain managers trace back the cause of the damage in the supply chain.
-It is important to note that there are cases that cannot be unambiguously assigned to one of those categories, even by expert annotators.
-However, a sub-goal of this work is to assess to which degree an automated classification is possible.
-It is worth mentioning that there are plenty of other relevant surface defects, reducing the quality of banana crops that are not included in this study, e.g., insect damage, speckles, sunburn, maturity stains, malformed fruits, broken necks, or chemical residues.
-Banana annotations
-To obtain relative defect sizes, we also annotated segmentation masks of banana fruits.
-We decided to only annotate bananas semantically on a pixel level, omitting instance information for two reasons: First, as images often contain multiple bunches that are partly cropped or occluded, it is challenging to differentiate instances.
-Second, our use case does not require us to identify instances, as we are not interested in counting single bananas or bunches.
-We would rather distinguish focused foreground bunches from those that may appear in the background and, therefore, annotate two types of pixel classes: “Foreground Banana” and “Background Banana”.
-All pixels not annotated as a banana of defect are assigned the “Background” class.
-We annotated banana masks using LabelStudio and Segment Anything point prompts (see Section 2.3 for more details).
+Therefore, stakeholders often do not have large annotated datasets readily available for their specific use case, defects, and cultivars that could be used to set up these models [6].
+We argue that this domain has a particular need for computer vision models that work in the low-data regime, where images, annotations, or both are limited.
+Recent computer vision research shows a trend toward foundation models that are pre-trained on large (often unlabeled) datasets, enabled by advances in self-supervised learning [2].
+From a practical view, these models promise to address low-data regimes as they are more easily adaptable to complex downstream tasks in a zero or few-shot manner [16].
+Most foundation models can be adapted to various tasks (e.g., classification, detection, segmentation, depth estimation) in a specific data modality (e.g., images, text, time series) or across modalities [e.g., 17, 29].
+Other foundation models are designed to excel in a certain task [15, 27] or domain problem [11, 22, 19] within a modality.
+In this work, we address the defect-based grading of fresh produce under constrained data and annotation availability, which occurs in many real-world scenarios, particularly in decentralized supply chains.
+Our end goal is to count and estimate the size of visible surface defects (i.e., surface blemishes such as bruises and scars) on fresh produce, which is essential for grading.
+We formulate this problem as a panoptic segmentation task [14].
+While our experiments focus on bananas as a representative commodity, the proposed methodology is generalizable to other types of fresh produce with comparable grading criteria.
+Our contribution is two-fold: First, we address the issue of label scarcity that naturally occurs with dense computer vision tasks such as semantic or panoptic segmentation.
+We will evaluate the Segment Anything Model (SAM) [15], a promptable foundation model for image segmentation, for the weakly supervised generation of panoptic masks.
+More specifically, we evaluate how defect instance masks align with dense human annotations and look for systematic failure cases.
+Second, using these predicted masks, we train a panoptic segmentation model on 476 images.
+We evaluate results against both human-annotated ground truth and masks generated by SAM.
+This allows us to estimate how the error introduced by using SAM-generated masks propagates through the downstream task.
+To our knowledge, this is one of the first works to combine foundation models like SAM with panoptic segmentation for defect quantification in the agricultural domain.
+2 Methods
 Figure 1: Overview of our approach.
@@
 This enables us to determine the number and size of visible surface defects from photographs of banana fruits.
-We validate our approach using a dataset of 476 images.
-2.2 Image segmentation
-Image segmentation uses computer vision techniques to analyze and understand images on a pixel level.
-In this section, we describe different paradigms of image segmentation and their corresponding evaluation metrics that we use to evaluate the quality of our models.
-2.2.1 Semantic Segmentation
-“Semantic segmentation” is concerned with classifying each pixel of an image into a predefined category (e.g., background, banana fruit, defect) without distinguishing between individual objects (instances) of the same class.
-(Ronneberger et al., 2015; Chen et al., 2017; Xie et al., 2021)
-Semantic segmentation maps are evaluated against a groundtruth using the Intersection-over-Union (IoU) score for a single class or the mean IoU (mIoU) score for all $N$ classes in a dataset (Everingham et al., 2012; Lin et al., 2014).
-Eqs.
-1 and 2 show the definition of those metrics where $G$ refers to the set of groundtruth pixels that are assigned to a class and $P$ to the set of prediction pixels accordingly.
-$IoU(G,P)$ $=\frac{|G\cap P|}{|G\cup P|}$ (1)
-$mIoU$ $=\frac{1}{N}\sum_{i=1}^{N}IoU_{i}$ (2)
-2.2.2 Instance Segmentation
-Instance segmentation goes a step further by categorizing each pixel and identifying each object instance of the same class separately.
-This is realized by assigning each pixel to one of potentially multiple objects and to a semantic class simultaneously.
-This technique is essential for tasks requiring counting or differentiating between items.
-It is considered more complex, combining object detection with pixel classification.
-(He et al., 2017; Lin et al., 2017)
-In terms of evaluation, instance segmentation is equivalent to object detection, which in machine learning describes the localization of instances via rectangular bounding boxes.
-The only difference is that the overlap of instances is defined by masks rather than boxes.
-There are several widely recognized metrics that compare a set of groundtruth bounding boxes or instance masks to a set of predicted bounding boxes or instance masks.
-Generally, a predicted and an annotated instance are matched if their localization exceeds a certain IoU threshold.
-As a first step, we calculate the precision $P$ and recall $R$ as follows:
-$P$ $=\dfrac{TP}{TP+FP}$ (3)
-$R$ $=\dfrac{TP}{TP+FN}$ (4)
-where $TP$ is the number of true positives, $FP$ is the number of false positives, and $FN$ is the number of false negatives.
-By definition, precision and recall are inversely related but must concur effectively to ensure that both relevant and correct instances are identified.
-The primary evaluation metric for object detection is Average Precision ($AP$) for a single class and mean Average Precision ($mAP$) for multiple classes.
-We follow established practices and report three different variations:
-  • $AP@50$: $AP$ at a $IoU$ threshold of $0.50$ as the most common metric in many benchmarks (e.g., Everingham et al., 2012),
-  • $AP@75$: $AP$ at a $IoU$ threshold of $0.75$ as a more strict metric,
-  • $AP$ (or $AP^{IoU=.50:.05:.95}$): $AP$ averaged over ten equidistant $IoU$ thresholds between $0.50$ and $0.95$ as proposed by Lin et al. (2014).
-The average recall is another evaluation metric used to measure the assertiveness of object detectors for a given class (Padilla et al., 2021).
-Similar to $AP$, the $AR$ metric evaluates a range of IOU thresholds.
-We use the equation provided by Hosang et al.
-(2015) to calculate $AR$ from a discrete set of thresholds based on the total number of ground truth instances $G$ in an image:
-$AR=\dfrac{2}{G}\sum_{i=1}^{G}max(IoU_{i}-0.5,0)$ (5)
-2.2.3 Panoptic Segmentation
-In scenarios where one requires both instance-level information for some classes that are countable (widely referred to as “things”) and semantic information about other non-countable components (referred to as “stuff”), these two concepts need to be combined into an approach called “Panoptic Segmentation”.
-This approach seeks to overcome the limitations of its predecessors by delivering a comprehensive segmentation map that includes both class labels and instance identification.
-(Cheng et al., 2021, 2022; Kirillov et al., 2019a, b)
-In our study, we declare defects as “things” and bananas as “stuff” as we are not interested in distinguishing single banana fruits. Thus, we simplify the machine learning task.
-Kirillov et al.
-(2019b) propose Panoptic Quality ($PQ$) as a comprehensive metric for panoptic image segmentation (Eq.
-6).
-The intuition behind $PQ$ is to provide a unified measure that captures the accuracy of detecting and correctly segmenting individual objects (defects in our case) as well as the overall segmentation quality of the entire scene.
-$PQ$ combines precision (how many predicted segments are correct) and recall (how many true segments are detected) into a single score, rewarding models that can accurately identify object boundaries and classify them correctly.
-$PQ=\frac{\sum_{(p,g)\in TP}\text{IoU}(p,g)}{|TP|+\frac{1}{2}|FP|+\frac{1}{2}|FN|}$ (6)
-In this work, we use Panoptic Quality as the primary evaluation metric to compare how well trained image segmentation models perform.
-2.3 Segment Anything Model (SAM)
-The Segment Anything Model (SAM) is a recently published foundation model for image segmentation (Kirillov et al., 2023).
+We validate our approach using a dataset of 476 images and 1440 annotated defects.
+2.1 Framing the problem
+Our work focuses on a common form of visual inspection relevant across a wide range of commodities: detecting surface defects, such as blemishes resulting from external impacts [1].
+Although grading criteria vary by commodity and market, we identify two key indicators that are consistently important: (a) the number of visible blemishes and (b) the size of each of those.
+Additionally, the type of defect, while secondary, provides valuable insight into potential causes, aiding supply chain traceability and intervention.
+We argue that an image segmentation approach is better suited for this task than regression or ordinal classification.
+Segmentation outputs are inherently interpretable and can be visually validated, which is crucial in quality assurance settings where transparency and accuracy are essential.
+To estimate the number of defects, it is essential to distinguish between individual defect instances.
+Additionally, identifying banana fruits is necessary for relative size estimation (see Section 2.6 for details); however, distinguishing between different banana fruit instances is not required beyond distinguishing foreground and background bunches in an image.
+Based on these requirements, we conclude that the panoptic segmentation [14] framework—combining both semantic and instance segmentation—is well-suited to our task.
+In principle, a standard semantic segmentation approach combined with postprocessing could also be used to separate instances (see Section 2.5), but panoptic segmentation offers a more integrated and efficient solution.
+2.2 Data collection
+We collected 476 images of single banana bunches containing at least one visible bruise or scar in three different sorting and packing facilities in India, located in Tembhurni (Maharashtra state), Raver (Maharashtra state), and Kamrej (Gujarat state).
+The images were taken with different smartphone cameras between harvesting and packing.
+To ensure independent data samples, each bunch was photographed only once.
+The photos have realistic, non-standardized properties: the proximity to the fruits, lighting conditions, image sharpness, and resolution can vary.
+Also, there can be other bananas and objects occasionally in the background.
+Such conditions are typical for images taken in supply chains by inspectors, so we wanted our method to accommodate that.
+We aim to first automatically detect, segment, and classify defects using a machine learning model.
+In the second step, we derive the number of visible surface defects and their relative size from these predictions.
+Defect annotations
+The images were annotated by a single expert using the open-source tool “LabelStudio” [24].
+As we are interested in the number of defects, we annotated defects on an instance level, meaning that we annotated one bounding box per defect.
+While a key property of our approach is to only use bounding boxes and generate instance masks via SAM, we also annotate defect masks for validation purposes.
+Specifically, each defect was one of four categories: “Old Bruise”, “New Bruise”, “Old Scar”, and “New Scar”.
+Bruises typically result from blunt force or pressure, which can arise from various factors such as handling, packing, impacts, or weight.
+Scars, on the other hand, are caused by sharp contact due to reasons like friction from leaves or stems, cuts from knives, crates, rough handling, laborers’ nails, or the sharp tips and crowns of neighboring bunches.
+Old defects are usually darker than new defects as they oxidize over time (see Figure S.1).
+In total, 1440 defects were annotated.
+(37 Old Bruise, 182 New Bruise, 387 Old Scar, 834 New Scar) While the automated detection of generic defects without classification is already an impactful use case, we chose to add this categorization as it helps supply chain managers trace back the cause of the damage in the supply chain.
+It is important to note that there are cases that cannot be unambiguously assigned to one of those categories, even by expert annotators.
+However, a sub-goal of this work is to assess to which degree an automated classification is possible.
+It is worth mentioning that there are plenty of other relevant surface defects reducing the quality of banana crops that are not included in this study, e.g., insect damage, speckles, sunburn, maturity stains, malformed fruits, broken necks, or chemical residues.
+Banana annotations
+To obtain relative defect sizes, we also annotated segmentation masks of banana fruits.
+We decided to annotate bananas semantically only on a pixel level, omitting instance information for two reasons: First, as images often contain multiple bunches that are partly cropped or occluded, it is challenging to differentiate instances.
+Second, our use case does not require us to identify instances, as we are not interested in counting single bananas or bunches.
+We would rather distinguish focused foreground bunches from those that may appear in the background and, therefore, annotate two types of pixel classes: “Foreground Banana” and “Background Banana”.
+All pixels not annotated as a banana of defect are assigned the “Background” class.
+We annotated banana masks using LabelStudio and Segment Anything point prompts (see Section 2.3 for more details).
+2.3 Mask generation with SAM
+The Segment Anything Model (SAM) is a recently published foundation model for image segmentation [15].
 It was trained on a large-scale dataset comprising diverse images from various domains, including natural scenes, urban environments, medical imagery, and other contexts.
@@
 Figure 1 (left) shows a high-level overview of SAM’s components.
-The model consists of three major components: (1) a vision transformer (Dosovitskiy et al., 2020) pre-trained with the Masked Autoencoder method (He et al., 2021) that serves as an image encoder; (2) a prompt encoder that allows the user to specify bounding boxes or points to localize the region of interest in an image [^1] In their paper, the authors of SAM mention open vocabulary text as input prompts.
-However, they did not publish their code for this feature as the results are not as good as for other prompts (REF needed).
-Therefore, we do not consider this option in this research.; (3) a mask decoder that generates segmentation masks based on the concatenated image and prompt embeddings.
+The model consists of three major components: (1) a vision transformer [7] pre-trained with the Masked Autoencoder method [9] that serves as an image encoder; (2) a prompt encoder that allows the user to specify bounding boxes or points to localize the region of interest in an image; (3) a mask decoder that generates segmentation masks based on the concatenated image and prompt embeddings.
 This approach yields impressive results for zero-shot instance segmentation in various domains.
-However, compared to conventional segmentation models, it comes with two disadvantages that require further adaptations or post-processing to enable SAM for a typical image segmentation use case: Firstly, SAM requires weak labels such as bounding boxes or point coordinates as input prompts.
-Secondly, the output masks are uncategorized, which hinders zero-shot usage of SAM in a multi-class setting.
-SAM2 (Ravi et al., 2024) is a recent extension of SAM to the video domain.
-Although videos are not the focus of this work, SAM2 has demonstrated that their video-based pre-training enhances zero-shot mask generation for images as well.
-In our study, we assess both SAM and SAM2 using a ViT-L backbone for zero-shot mask generation of banana fruits and defects.
-2.4 Experimental setup
-Figure 2 shows an overview of our approach.
-The process starts with an image dataset with coarsely annotated labels: bounding boxes to locate and assign labels to defects and sets of points that identify foreground and background bananas.
-We utilize a promptable foundation model for image segmentation (SAM) to generate fine-grained pixel-wise annotations of each image in the dataset based on the coarse hand annotations.
-It is important to note that SAM is not trained in this procedure but is used here in a zero-shot manner to infer dense annotations only.
-These newly created annotations are used to train a panoptic segmentation model in a standard supervised machine learning workflow.
-The panoptic segmentation model we use is a Maskformer (Cheng et al., 2021) model pre-trained on the ADE20k dataset (Zhou et al., 2017).
-Further details on model training, such as hyperparameters and data augmentation, can be found in A.
+However, compared to conventional segmentation models, it comes with two disadvantages that require further adaptations or post-processing to enable SAM for a typical image segmentation use case: Firstly, SAM requires weak labels, such as bounding boxes or point coordinates, as input prompts.
+Secondly, the output masks are generally uncategorized.
+We explore the use of SAM in a zero-shot setting to generate dense panoptic masks from coarse annotations—specifically, bounding boxes for defects and point annotations for banana fruits—effectively serving as a semi-automated annotation tool.
+However, it is not immediately clear whether SAM-generated masks sufficiently align with human-annotated masks.
+Other than common objects in images, surface defects potentially have ambiguous boundaries that are subject to expert judgment.
+Validating SAM masks
+We validate SAM masks in two ways: First, we directly evaluate the alignment between SAM-generated and human-annotated masks under different conditions, including ablations on the backbone model size and the resolution of the defect areas in the image (Section 3.1).
+Second, we assess the performance of our panoptic segmentation model against both human-annotated and SAM-generated masks (Section 3.2).
+This allows us to draw conclusions about the extent of error propagation resulting from inaccurate mask generation.
+More specifically, we report panoptic evaluation using three different combinations of mask sources for training and validation data: (a) both training and validation use hand-annotated defect masks; (b) both training and validation use SAM-generated defect masks; and (c) training uses SAM-generated masks while validation uses hand-annotated masks.
+By comparing the evaluation metrics across these three variations, we can quantify the trade-off between the time-saving benefits of automated mask generation and the quality of the resulting output.
+2.4 Image segmentation
+In Section 2.1, we argued why we frame this use case as a hand as a panoptic segmentation problem.
+This approach yields a dense semantic segmentation mask for the whole image while instance identification is provided for certain specified categories only [4, 5, 13, 14].
+In our study, we declare defects as “things” (countable) and bananas as “stuff” (non-countable) as we are not interested in distinguishing single banana fruits.
+We train a Maskformer [4] pre-trained on the ADE20k dataset [30] in a standard supervised learning workflow (Figure 1, right).
+Evaluation metrics
+We evaluate the final model using metrics from all three paradigms: semantic, instance, and panoptic segmentation.
+Semantic segmentation maps are evaluated against a groundtruth using the Intersection-over-Union (IoU) score for a single class or the mean IoU (mIoU) score for all classes in a dataset [8, 18].
+For instance detections, we follow the established practice to report average precision at different $IoU$ threshold ($AP@50$, $AP@75$) [8].
+$AP$ is equivalent to $AP^{IoU=.50:.05:.95}$, averaged over ten equidistant $IoU$ thresholds between $0.50$ and $0.95$ as proposed by Lin et al.
+[18].
+For the average recall metric ($AR$), we follow the definition of Hosang et al.
+[10] evaluating a range of IOU thresholds.
+In addition, we use Panoptic Quality ($PQ$) as introduced by Kirillov et al.
+[14] as an integrated metric for panoptic segmentation.
 Cross-validation
@@
 The evaluation metrics are then reported as the mean and standard deviation across these five experiments.
-Validating SAM masks.
-Qualitative analysis indicated that SAM-generated defect masks are generally plausible.
-However, to quantify the impact of training on SAM-generated masks compared to using hand-annotated pixel-wise masks, we repeat our experiments with different combinations of mask sources for training and validation data: (a) both training and validation use hand-annotated defect masks; (b) both training and validation use SAM-generated defect masks; and (c) training uses SAM-generated masks while validation uses hand-annotated masks.
-By comparing the evaluation metrics across these three variations, we can quantify the trade-off between the time-saving benefits of automated mask generation and the quality of the resulting output.
+Further details on model training, such as hyperparameters and data augmentation, can be found in Appendix B.
 2.5 Postprocessing
 In some cases, it can be ambiguous whether to treat closely located defects as a single defect or multiple defects.
-To address this and ensure more consistent defect separation, we introduce a simple algorithm-based postprocessing method with the following three steps: (1) Combine all defect instances of the same type into a single binary map.
-(2) Use a connected components algorithm (OpenCV, 2015) to generate new instances where no pixels are connected.
-3) Apply a dilation operation to expand all instance masks by $d$ pixels (we choose $d=5$) in all directions.
+To address this and ensure more consistent defect separation, we introduce a simple algorithm-based postprocessing method with the following three steps:
+  1. 1.
+    Combine all defect instances of the same type into a single binary map (skip if a semantic segmentation model is used).
+  2. 2.
+    Use a connected components algorithm [20] to generate new instances where no pixels are connected.
+  3. 3.
+    Apply a dilation operation to expand all instance masks by $d$ pixels (we choose $d=5$) in all directions.
 This expanded auxiliary mask is then used to identify overlapping instances, which are subsequently merged.
 In other words, defect instances of the same class within a distance of $2d$ pixels are merged.
-Figure 3: Distribution of IoU values when comparing annotated and SAM-generated defect masks. We generally observe a high overlap, while SAM2 masks are closer aligned to the hand-annotated ones.
+Note that after step 1, we have regular semantic segmentation representation, and thus, using steps 2 and 3, one could also obtain panoptic results based on standard semantic segmentation outputs.
+However, during initial experimentation, we found that training a Maskformer [4] yields better results than a comparable semantic segmentation model (Segformer [26]).
+Also, we consider this postprocessing as optional when a panoptic model is used.
 2.6 Deriving number and size of defects
-In the context of quality estimation, we are interested in two practical indicators that can be derived from an image’s panoptic prediction: The number and size of defects.
-While the first can be estimated by counting the number of defect instances, an accurate metrical size estimation in a two-dimensional image is impossible due to unknown scale and distortions.
+For many real-world grading applications, one is interested in the number and size of defects, which can both be derived (or approximated) from panoptic masks.
+While the first can be estimated by counting the number of defect instances, an accurate metric size estimation from a single image is generally impossible due to unknown scale, perspective distortion, and lack of depth information.
 Thus, we use the relative defect size, simplified as the number of visible defect pixels divided by the number of visible foreground banana pixels, as a proxy.
-3 Results and Discussion
-3.1 Comparison of annotated and generated masks
-Our approach uses SAM to generate dense segmentation masks in a zero-shot setting (i.e., without pre-training).
-To assess the quality of these masks, we first compare the masks produced by different SAM models (SAM and SAM2) with hand-annotated ground truth masks (Figure 3).
-Overall, we observe a high overlap between the hand-annotated defect masks and those generated by both SAM models.
-Notably, SAM2 exhibits even greater accuracy than SAM, with 90.5% of its masks achieving an IoU score of 0.5 or higher.
-Based on these findings, we conclude—and later demonstrate—that our zero-shot mask generation approach is both valid and capable of producing segmentation results comparable to hand-annotated masks.
-Figure 4: Hand-annotated vs.
-estimated defect counts per image.
-We predict the exact number of defects 36.5% of the time (green squares).
-78.3% of the time, we predict correctly within a ±$\pm 1$ tolerance (orange squares).
-Generally, our model tends to predict more defects in the images compared to the annotations.
-Figure 5: Annotated vs estimated defect sizes.
-We pairwise match those defect instances with the highest IoU agreement between annotation and estimation (minimum 0.5) and calculate their sizes relative to the corresponding foreground banana masks.
-Relative size prediction is generally accurate with a Pearson correlation of $r=0.96,n=829$.
-Each blue dot is one pair of defects, the orange line is a fitted regression line, black line is $x=y$ diagonal.
-3.2 Evaluation with image segmentation metrics
+3 Results
+3.1 SAM defect masks
+Figure 2: Agreement between human-annotated and SAM-predicted masks by defect size (number of pixels).
+The x-axis shows binned size categories for defect sizes in pixels (per annotated mask), while $n$ denotes the number of samples in each bin.
+The y-axis shows the agreement between annotated and predicted masks (IoU).
+SAM fails to align with human annotations for small ($<\!100$ pixels) and very small ($<\!10$ pixels) defects.
+To understand the failure cases of larger defects ($>\!100$ pixels), we added exemplary visualizations in Figure 3.
+Those cases often show thin, long scars.
+We can also see that a larger SAM model size (ViT-H) leads to consistently higher agreement with human annotations across all defect sizes.
+An additional analysis of the impact of SAM model sizes can be found in Figure S.2.
+Figure 3: SAM failure cases.
+Exemplary examples where SAM (ViT-L) fails to align with the annotated masks.
+We picked examples with the lowest IoU where the annotated mask size is $>\!100$ pixels.
+The examples shown illustrate that long, thin scars are difficult to capture.
+Only one defect per image is shown.
+How well do SAM-predicted masks align with human annotations?
+We use SAM in a zero-shot manner without any fine-tuning.
+Thus, it is not self-evident that all defects with different shapes and sizes align closely with human annotations.
+Also, the choice of the SAM backbone might have an influence on predictions as model capacity usually correlates with model accuracy.
+Figure 3 compares the agreement between hand-annotated and SAM-generated masks for two different SAM backbone sizes (ViT-L and ViT-H) and for different defect sizes defined by the number of pixels in the annotated mask.
+First, we can see that a larger SAM model size (ViT-H) leads to consistently higher agreement with human annotations across all defect sizes.
+This finding is confirmed by an additional analysis where we plot the IoU distributions for three different SAM backbones (Figure S.2).
+However, we see that the alignment is generally high, and failure cases are long-tail outliers for all model backbones.
+A second observation from Figure 3 is that SAM fails to align with human annotations for small ($<\!100$ pixels) and very small ($<\!10$ pixels) defects.
+To understand the failure cases of larger defects ($>\!100$ pixels), we added exemplary visualizations in Figure 3.
+In those cases, which often show long, thin scars, SAM produces masks that are too large.
+3.2 Panoptic predictions
+Table 1: Image segmentation metrics.
+All results are reported via their mean and standard deviation over the five different cross-validation splits.
+We compare different setups where target defect masks are either annotated (“Anno.”) or generated by SAM (-L/-H refers to the size of the ViT backbone).
+Expectedly, when Maskformer is trained on SAM-generated masks, the defect IoU is slightly lower when evaluating annotated masks.
+In the same comparison, recall and precision are degrading less strongly.
+Postprocessing (“PP”) significantly improves precision and recall of defects.
+Using SAM-generated instead of annotated training masks, leads to identical PQ ($\mathord{\sim}0.779$), while mIoU is slightly worse ($\mathord{\sim}0.808$ vs $\mathord{\sim}0.810$).
+A larger SAM backbone size does not lead to an improvement in downstream segmentation quality.
+The configuration highlighted in blue is used for subsequent visualizations.
+Defect masks Defects Fruits Overall Model PP train val AP AP@50 AP@75 AR IoU IoU FG IoU BG mIoU PQ Maskformer Anno.
+Anno.
+.$153\pm.040$$263\pm.066$$163\pm.050$$262\pm.051$$563\pm.037$$956\pm.006$$750\pm.050$$810\pm.018$$764\pm.010$ Maskformer Anno.
+Anno.
+.$223\pm.030$$413\pm.059$$2$040 .$397\pm.034$$563\pm.037$$956\pm.006$$750\pm.050$$810\pm.018$$779\pm.010$ Maskformer SAM-L Anno.
+.$129\pm.011$$231\pm.019$$134\pm.017$$243\pm.024$$559\pm.024$$952\pm.007$$746\pm.056$$808\pm.018$$760\pm.008$ Maskformer SAM-L Anno.
+.$217\pm.020$$405\pm.038$$211\pm.024$$393\pm.035$$562\pm.028$$952\pm.007$$746\pm.056$$808\pm.018$$779\pm.008$ Maskformer SAM-L SAM-L .$137\pm.015$$232\pm.015$$147\pm.023$$253\pm.025$$574\pm.022$$952\pm.007$$746\pm.056$$811\pm.018$$762\pm.006$ Maskformer SAM-L SAM-L .$231\pm.023$$409\pm.050$$235\pm.026$$410\pm.039$$576\pm.025$$952\pm.007$$746\pm.056$$812\pm.018$$782\pm.007$ Maskformer SAM-H Anno.
+.$150\pm.016$$257\pm.027$$160\pm.021$$251\pm.033$$547\pm.063$$952\pm.008$$733\pm.045$$801\pm.023$$759\pm.007$ Maskformer SAM-H Anno.
+.222±.028plus-or-minus.222.028.$2$028 .$418\pm.056$$211\pm.031$$384\pm.048$$548\pm.064$$952\pm.008$$733\pm.045$$801\pm.023$$7$010 Maskformer SAM-H SAM-H .$158\pm.022$$255\pm.024$$175\pm.031$$258\pm.033$$573\pm.067$$952\pm.008$$733\pm.045$$807\pm.024$$763\pm.007$ Maskformer SAM-H SAM-H .$242\pm.031$$422\pm.048$$251\pm.043$$403\pm.058$$574\pm.068$$952\pm.008$$733\pm.045$$808\pm.024$$783\pm.011$
 Next, we assess image segmentation results on validation images using the metrics outlined in Section 2.4.
-Table 1 presents these metrics across various experimental setups: In the first experiment, we used hand-annotated defect masks for both training and validation, achieving an average panoptic quality (PQ) of 77.5% after applying postprocessing.
+Table 1 presents these metrics across various experimental setups: In the first experiment, we used hand-annotated defect masks for both training and validation, achieving an average panoptic quality (PQ) of 77.9% after applying postprocessing.
 This serves as our baseline for assessing how segmentation quality changes when replacing hand-annotated masks with SAM-generated ones.
 In the second experiment, we employ SAM-generated masks for both training and validation.
-The results show comparable PQ scores of 77.3% (SAM) and 77.6% (SAM2).
+The results show a slightly higher PQ score of 78.2% (ViT-L backbone).
 In the final experiment, SAM-generated masks are used for training, while hand-annotated masks are used for validation.
-Here, the PQ scores are approximately $\mathord{\sim}0.5\%$ lower for both SAM and SAM2 compared to when hand-annotated masks are used for training.
-In conclusion, while SAM-generated masks are not fully equivalent to hand-annotated ones—occasionally failing to capture certain edge cases—they still produce metrics that are sufficiently close to be considered a viable low-effort alternative: For our dataset, we estimated an annotation time of approx.
-15 seconds per image for coarse annotations and 3-5 minutes per image for pixel-wise annotations.
-An additional interpretation of these results is that there is a slight but consistent difference in how our human annotator and SAM delineate defect boundaries.
-Nonetheless, the Maskformer model demonstrates the ability to learn from both annotation styles effectively.
-Lastly, it is evident that our postprocessing method is effective for improving defect masks.
-3.3 Qualitative evaluation
-The results can be qualitatively analyzed through the example visualizations in Figure 6 showing the segmentation using a single generic defect class.
+One can see a slight drop in detection metrics (AP: $-0.6\%$, AR: $-0.4\%$, IoU: $-0.1\%$), while the overall PQ score is almost identical when SAM-generated instead of hand-annotated masks are used for training.
+While we previously showed that the larger SAM variant (ViT-H) has a higher defect mask agreement with hand annotations (see Figure 3), this does not, somewhat counterintuitively, lead to better segmentation metrics.
+This suggests that the downstream model (MaskFormer, in our case) can compensate for a certain degree of inaccuracy in the training masks.
+Across all settings, IoU scores for segmenting foreground bananas are consistently high ($>95\%$), while those for background bananas are noticeably lower ($\mathord{\sim}74\%$).
+This suggests that the model rarely confuses the two classes—if it did, both scores would be low.
+Instead, we conclude that background bananas are occasionally missed in the annotations.
+Qualitative comparisons support this interpretation (see, for example, Figure 6, second and fourth row).
+Furthermore, the postprocessing described in Section 2.5 yields a significant improvement in AP across all settings.
+The panoptic results can be qualitatively analyzed through the example visualizations in Figure 6 showing the segmentation using a single generic defect class.
 In general, the masks appear sharp, particularly for bananas, and the model reliably distinguishes between fruits in the foreground or center of the image and those in the background.
@@
 This underscores the challenge of defining ground-truth defect annotations in these images.
+Figure 4: Annotated versus predicted defect counts per image.
+We predict the exact number of defects 36.2% of the time (green squares).
+76.2% of the time, we predict correctly within a ±$\pm 1$ tolerance (orange squares).
+Generally, our model tends to predict more defects in the images compared to the annotations.
+Figure 5: Annotated versus predicted defect sizes.
+We pairwise match those defect instances with the highest IoU agreement between annotation and estimation (minimum 0.5) and calculate their sizes relative to the corresponding foreground banana masks (non-matchable defects are excluded in this analysis).
+Relative size prediction is generally accurate with a Pearson correlation of $r=0.96,n=793$.
+Each blue dot is one pair of defects, the orange line is a fitted regression line, and the black dashed line is the $x=y$ diagonal.
+Lastly, we also conducted experiments using four distinct defect types instead of a single defect class.
+We found that defect categorization is very unreliable and attribute this mainly to ambiguous annotations.
+Thus, we do not recommend using our dataset for defect categorization at this point (see Appendix D for a more detailed discussion).
+Figure 6: Example visualizations of annotated vs predicted masks.
+Left: Input Image, Mid: Annotation, Right: Maskformer Prediction.
+Red rectangles enclose defect instances.
+Segments are color-coded as follows: Foreground Banana, Background Banana, Defect,
 3.3 Number and size of defects
@@
 We compare the number of defects detected per image (Figure 4) and the relative defect sizes (Figure 5) by our best model (x-axes) and the human annotations (y-axes).
-The model accurately matches the exact number of defects in 36.5% of cases and is off by one in 78.3% of cases. This further highlights the issue of ambiguous ground truth, as discussed earlier.
+The model matches the exact number of defects in 36.2% of cases and is off by one in 76.2% of cases.
 For relative size analysis, we pairwise match those defect instances with the highest IoU agreement between annotation and estimation.
@@
 annotated relative sizes ($r=0.96)$, while the model tends to slightly underestimate sizes, visible by the different slopes of the regression line (orange) and $x=y$ (black dashed) lines in Figure 5.
-Table 1: Main results with a single defect category.
-All results are reported via their mean and standard deviation over the five different cross-validation splits.
-We compare different setups where target defect masks are either annotated (“Anno.”) or generated by SAM/SAM2.
-Expectedly, when Maskformer is trained on SAM-generated masks, the defect IoU is slightly lower when evaluating annotated masks.
-However, recall and precision are comparable to the fully annotated setting.
-We also demonstrate that our postprocessing (“PP”) significantly improves precision and recall of defects.
-Banana segmentation quality is slightly better in the fully annotated setting ($\mathord{\sim}0.81$ vs $\mathord{\sim}0.80$ mIoU).
-The configuration highlighted in blue is used for subsequent visualizations.
-Defect masks Defects Fruits Overall Model PP train val AP AP@50 AP@75 AR IoU IoU FG IoU BG mIoU PQ Maskformer Anno.
-Anno.
-.$133\pm.022$$243\pm.027$$134\pm.027$$256\pm.036$$593\pm.032$$952\pm.005$$732\pm.034$$812\pm.010$$758\pm.010$ Maskformer Anno.
-Anno.
-.$220\pm.019$$422\pm.014$$215\pm.033$$412\pm.041$$593\pm.032$$952\pm.005$$732\pm.034$$813\pm.010$$775\pm.007$ Maskformer SAM Anno.
-.$128\pm.030$$234\pm.043$$128\pm.040$$238\pm.025$$546\pm.040$$951\pm.006$$731\pm.052$$800\pm.011$$752\pm.011$ Maskformer SAM Anno.
-.$205\pm.017$$392\pm.040$$197\pm.013$$380\pm.022$$548\pm.039$$951\pm.006$$731\pm.052$$801\pm.011$$770\pm.005$ Maskformer SAM SAM .$137\pm.030$$241\pm.045$$4.1$034 .$250\pm.027$$561\pm.038$$951\pm.006$$731\pm.052$$804\pm.012$$755\pm.010$ Maskformer SAM SAM .$220\pm.016$$404\pm.044$$2$016 .$399\pm.023$$562\pm.037$$951\pm.006$$731\pm.052$$804\pm.011$$773\pm.005$ Maskformer SAM2 Anno.
-.$121\pm.028$$227\pm.047$$121\pm.036$$238\pm.034$$5.5$025 .$950\pm.005$$720\pm.041$$799\pm.015$$753\pm.011$ Maskformer SAM2 Anno.
-.$201\pm.022$$395\pm.018$$184\pm.030$$392\pm.020$$554\pm.025$$950\pm.005$$720\pm.041$$799\pm.015$$770\pm.013$ Maskformer SAM2 SAM2 .$134\pm.039$$229\pm.050$$139\pm.054$$249\pm.044$$579\pm.028$$950\pm.005$$720\pm.041$$805\pm.015$$757\pm.012$ Maskformer SAM2 SAM2 .$225\pm.035$$405\pm.029$$221\pm.056$$415\pm.034$$579\pm.028$$950\pm.005$$720\pm.041$$805\pm.015$$776\pm.012$
-3.5 Defect Categorization
-We also conducted experiments using four distinct defect types instead of a single defect class (see B for details).
-In this setup, the model must not only detect and segment defects but also categorize them into one of four pre-defined classes.
-Our results clearly show that while the detection and segmentation of bruises and scars work well, a reliable categorization into one of the four predefined classes is not possible with the current approach and dataset.
-We hypothesize that this limitation stems from one or more of the following factors:
-  1. 1.
-    Ambiguous annotation: The classification of defects across the four predefined categories may be subjective to a large degree.
-The distinction between “old” and ”new“ defects is not always clear-cut, as the transition is gradual.
-Additionally, distinguishing bruises from scars can be ambiguous, especially for non-experts (see Figure 1).
-  2. 2.
-    Image resolution: We use $1024^{2}$ pixel resolution for images.
-While this is generally considered high for machine learning tasks, defects can be small and, thus, only be represented by a few pixels, making them harder to categorize.
-  3. 3.
-    The four defect types are unevenly represented in our dataset (37/82/387/834).
-A more balanced distribution of categories and a larger number of defect samples are likely to improve categorization accuracy.
-We recommend collecting more examples from the two underrepresented classes for future work.
-Figure 6: Example visualizations of annotated vs predicted masks using a single defect category.
-Left: Input Image, Mid: Annotation (SAM), Right: Maskformer Prediction.
-Segments are color-coded as follows: Foreground Banana, Background Banana, Defect,
-4 Conclusion and Outlook
-In this work, our objective was to have an efficient way to implement machine learning to count the number of defects on banana fruit and have their relative sizes.
-We demonstrated the use of a panoptic image segmentation model to tackle the detection and pixel-wise segmentation of bananas and surface blemishes in images.
-Our approach significantly reduces the effort required for hand annotation by relying on coarse labels.
-Specifically, we utilize a pre-trained foundation model for image segmentation (the Segment Anything Model) to generate the dense annotations necessary for training the panoptic model.
-This method reduces the manual annotation workload by at least a factor of 10 while maintaining high-quality results: the measured Panoptic Quality drops from 77.5% to 77.0%) w.r.t.
-the annotated targets.
-Our results generally demonstrate reliable and plausible segmentation of both foreground and background bananas, as well as a single generic defect category.
-However, when compared to human annotations, the model matches the exact number of defects in only 36.5% of cases and is off by up to one defect in 78.3% of cases.
-While some of this discrepancy can be attributed to model performance, it is more likely due to the target annotations.
-Unlike for other machine learning applications where target labels are often close to a “ground truth”, the identification of significant blemishes is to a certain degree subjective.
-An improvement on the labeling side could be to employ multiple annotators and measure their level of agreement.
-The benefits of using (noisy) annotations from multiple annotators for samples that are non-trivial to label have been explored in the machine learning literature (Welinder et al., 2010; Sheng et al., 2008).
-Thus, we recommend that future work carries out a more sophisticated annotation process with multiple expert annotators.
-The disagreement between annotators could then serve as a baseline for evaluating the performance of future machine learning models.
-In addition, we demonstrated that with our current method and dataset, it is not possible to reliably categorize defects into one of the four predefined classes, namely new scars, old scars, new bruises, and old bruises.
-A variable ground truth is likely a contributing factor, along with the skewed class distribution and the potentially low image resolution for small defect sizes.
-Addressing these challenges will require further investigation in future work to improve the reliability of defect categorization.
-Our method demonstrates the potential for dense image segmentation in low-data scenarios and can be broadly applied to any commodity with countable, localizable defects.
-However, for defects that are non-countable—such as speckles, sunburn, or chemical residues—a different machine learning paradigm (such as classification) would be more applicable.
+4 Discussion and Limitations
+In this work, we lay the foundation for the automated, defect-based grading of fresh produce using limited samples and annotations.
+We present a carefully curated dataset comprising 476 real-world images of banana bunches and 1,440 annotated surface defects covering a wide range of sizes, shapes, and categories.
+Further, we highlight the trade-off when using SAM for generating dense image masks from coarse annotations.
+Specifically, we estimate that the manual annotation workload can be reduced by at least a factor of 10 while image segmentation metrics for the downstream panoptic segmentation task experience only a moderate decrease (Defect IoU 56.3%→56.2%), Defect AP 22.3%→21.7%, Defect AR 39.7%→39.3%).
+We train a panoptic segmentation model and propose both a postprocessing algorithm and a procedure to extract grading-relevant metrics—such as defect count and relative size—from the model’s predictions.
+While it can be questioned if the current model is sufficient for practical deployment, it shows promise.
+Compared to human annotations, the model predicts the exact number of defects in 36.2% of cases and is off by at most one defect in 76.2% of cases.
+Although part of this discrepancy can be attributed to model limitations, we identify several avenues for improvement for future work, which we outline in Section 5.
+Lastly, it is important to note that countable and sizable defects represent only a subset of relevant grading criteria.
+For bananas, other defect types—such as speckles, sunburn, chemical residues, or malformed shapes—may require fundamentally different approaches, such as classification or ordinal regression.
+5 Future Work
+Improving model performance
+We see several ways to improve model performance on our dataset:
+  • Model choice: The landscape of deep learning model architectures evolves rapidly.
+Our approach will likely benefit from better SAM variants and improved panoptic segmentation architectures.
+SAM 2 [21] and Mask2Former [5] are the logical next candidates to be evaluated in our framework.
+  • Fine-tune SAM: Instead of using SAM in a zero-shot manner, one could fine-tune the model on domain data to achieve better segmentation masks for domain-specific instances.
+This was already successfully done in different contexts [see e.g., 11].
+We hypothesize that this will help with the issues of thin, long scars that can be seen in Figure 3.
+  • Categorized segmentation with SAM: Instead of our current two-model setup, adapting SAM to directly support categorized semantic or panoptic segmentation in a zero- or few-shot manner would provide a more elegant solution.
+Future research will likely explore such extensions.
+  • As we have highlighted in Figure 3, SAM fails with very small defect sizes. We expect that small defect predictions will improve when increasing the used image resolution (we use 10242).
+Improving data set creation
+We also want to highlight the issue of ambiguity in expert annotations for defect segmentation and, in particular, categorization.
+We would like to encourage a broader discussion on this challenge.
+One potential improvement would be to involve multiple annotators and quantify their agreement.
+The value of using (noisy) annotations from multiple annotators, especially for non-trivial labeling tasks, has been well documented in the machine learning literature [25, 23].
+We, therefore, recommend that future work adopt more sophisticated annotation protocols involving multiple experts, using inter-annotator disagreement as a baseline for evaluating machine learning models.
 Acknowledgments
@@
 Supplementary Materials
-The source code of all experiments, as well as pre-trained models, can be accessed on GitHub: https://github.com/manuelknott/banana-defect-segmentation.
-Supplementary Materials
-Appendix A Implementation details
+Appendix A Dataset visualizations
+Figure S.1: Defect examples.
+Example images of banana surface defects we aim to detect in this study.
+Bruises usually result from a dull impact, while scars are caused by a sharp impact.
+Old defects are usually darker than new defects as they oxidize over time.
+Appendix B Implementation details
 All images were resized and padded to $1024^{2}$ pixels resolution.
 Training images were augmented with a 50% Random Horizontal Flip and with additional random color-based augmentations (see Table S.1) to increase model robustness.
-Table S.2 shows the hyperparamters used for training the Maskformer models.
+Table S.2 shows the hyperparameters used to train the Maskformer models.
 We evaluated the model every five epochs on the validation set and saved the model with the highest Panoptic Quality.
-All trainings were done on a single NVIDIA RTX A5000 GPU with 24GB memory.
+All models were trained on a single NVIDIA RTX A5000 GPU with 24GB of memory.
 Table S.1: Random color augmentations applied to training images. Values are uniformly sampled from the specified ranges.
@@
 Evaluation frequency every 5 epochs
-Optimizer Adam (Kingma, 2014)
+Optimizer Adam [12]
 Learning rate $5\times 10^{-5}$
 Learning rate schedule constant
-Appendix B Results using four defect categories
+Appendix C Additional SAM evaluations
+Figure S.2: Distribution of IoU values when comparing annotated and SAM-generated defect masks.
+While the ViT-B and ViT-L variants show a similar distribution, using the largest model size (ViT-H) leads to significantly higher overlap with hand-annotated masks.
+Appendix D Results using four defect categories
+In this section, we describe our results when using four different defect classes instead of a single joint category.
+Our results clearly show that while the detection and segmentation of bruises and scars work well, a reliable categorization into one of the four predefined classes is not possible with the current approach and dataset.
+We hypothesize that this limitation stems from one or more of the following factors:
+  1. 1.
+    Ambiguous annotation: The classification of defects across the four predefined categories may be subjective to a large degree.
+The distinction between “old” and ”new“ defects is not always clear-cut, as the transition is gradual.
+Additionally, distinguishing bruises from scars can be ambiguous, especially for non-experts (see Figure S.1).
+  2. 2.
+    Image resolution: We use $1024^{2}$ pixel resolution for images.
+While this is generally considered high for machine learning tasks, defects can be small and, thus, only be represented by a few pixels, making them harder to categorize.
+  3. 3.
+    The four defect types are unevenly represented in our dataset (37/182/387/834).
+A more balanced distribution of categories and a larger number of defect samples are likely to improve categorization accuracy.
+We recommend collecting more examples from the two underrepresented classes for future work.
 Table S.3: Results using multiple defect categories.
@@
 Anno.
-.$027\pm.060$$011\pm.024$$058\pm.024$$186\pm.107$$039\pm.035$$229\pm.048$$063\pm.026$$376\pm.077$$047\pm.020$$493\pm.018$$471\pm.013$ Maskformer Anno.
+.$031\pm.059$$028\pm.043$$043\pm.024$$157\pm.037$$036\pm.018$$225\pm.097$$050\pm.018$$316\pm.054$$040\pm.019$$482\pm.024$$471\pm.023$ Maskformer Anno.
 Anno.
-.$009\pm.020$$051\pm.061$$063\pm.013$$115\pm.067$$030\pm.021$$199\pm.062$$034\pm.013$$123\pm.061$$034\pm.009$$447\pm.028$$442\pm.021$ Maskformer SAM Anno.
-.$0.000\pm.0$000 .$0.000\pm.0$000 .$058\pm.031$$166\pm.078$$031\pm.019$$243\pm.038$$064\pm.035$$351\pm.067$$038\pm.014$$488\pm.012$$462\pm.008$ Maskformer SAM Anno.
-.$0.000\pm.0$000 .$003\pm.006$$071\pm.024$$079\pm.046$$024\pm.018$$178\pm.055$$4.0$014 .$122\pm.035$$034\pm.010$$433\pm.014$$440\pm.007$ Maskformer SAM SAM .$0.000\pm.0$000 .$0.000\pm.0$000 .$061\pm.034$$173\pm.081$$035\pm.020$$247\pm.042$$069\pm.036$$359\pm.067$$041\pm.014$$490\pm.013$$467\pm.007$ Maskformer SAM SAM .$0.000\pm.0$000 .000±.001plus-or-minus.000.001.000$\pm$.$0$001 .$077\pm.026$$083\pm.049$$029\pm.024$$180\pm.058$$043\pm.014$$125\pm.036$$037\pm.010$$434\pm.016$$442\pm.006$ Maskformer SAM2 Anno.
-.$0.000\pm.0$000 .$0.000\pm.0$000 .$047\pm.019$$144\pm.082$$039\pm.038$$235\pm.086$$062\pm.014$$394\pm.052$$037\pm.007$$490\pm.021$$467\pm.009$ Maskformer SAM2 Anno.
-.$0.000\pm.0$000 .$021\pm.031$$064\pm.033$$127\pm.083$$027\pm.019$$168\pm.064$$033\pm.012$$094\pm.038$$031\pm.007$$437\pm.017$$437\pm.017$ Maskformer SAM2 SAM2 .$0.000\pm.0$000 .$0.000\pm.0$000 .$049\pm.019$$152\pm.086$$055\pm.065$$242\pm.090$$067\pm.012$$405\pm.053$$043\pm.014$$494\pm.022$$472\pm.011$ Maskformer SAM2 SAM2 .$0.000\pm.0$000 .$020\pm.030$$067\pm.036$$133\pm.086$$038\pm.034$$170\pm.066$$036\pm.013$$098\pm.039$$035\pm.003$$439\pm.018$$442\pm.019$
+.$030\pm.059$$013\pm.018$$088\pm.039$$114\pm.035$$034\pm.016$$223\pm.083$$017\pm.009$$066\pm.021$$042\pm.022$$436\pm.018$$431\pm.015$ Maskformer SAM-L Anno.
+.$039\pm.055$$060\pm.080$$034\pm.016$$175\pm.074$$045\pm.034$$179\pm.022$$066\pm.018$$345\pm.046$$046\pm.012$$487\pm.018$$477\pm.013$ Maskformer SAM-L Anno.
+.$032\pm.049$$031\pm.038$$074\pm.017$$124\pm.051$$038\pm.021$$161\pm.009$$032\pm.014$$104\pm.039$$044\pm.015$$437\pm.018$$435\pm.019$ Maskformer SAM-L SAM-L .$037\pm.053$$063\pm.081$$036\pm.016$$178\pm.075$$045\pm.033$$182\pm.023$$071\pm.019$$353\pm.049$$047\pm.012$$489\pm.018$$478\pm.013$ Maskformer SAM-L SAM-L .$032\pm.049$$027\pm.034$$8.0$018 .$127\pm.051$$036\pm.020$$163\pm.010$$034\pm.015$$108\pm.041$$046\pm.015$$437\pm.018$$436\pm.019$ Maskformer SAM-H Anno.
+.$019\pm.026$$074\pm.078$$043\pm.020$$157\pm.092$$031\pm.021$$252\pm.063$$056\pm.016$$318\pm.047$$037\pm.014$$495\pm.010$$474\pm.010$ Maskformer SAM-H Anno.
+.$032\pm.056$$030\pm.029$$093\pm.022$$145\pm.074$$025\pm.029$$230\pm.041$$029\pm.011$100±.050plus-or-minus.100.050.100$\pm$.$0$050 .$045\pm.016$$453\pm.013$$449\pm.009$ Maskformer SAM-H SAM-H .$025\pm.035$$092\pm.105$$049\pm.022$$163\pm.091$$037\pm.024$$264\pm.064$$062\pm.018$$328\pm.047$$043\pm.015$$501\pm.010$$482\pm.011$ Maskformer SAM-H SAM-H .$045\pm.078$$029\pm.027$$106\pm.022$$149\pm.073$$027\pm.028$$239\pm.041$$034\pm.010$$104\pm.051$$053\pm.019$$455\pm.013$$456\pm.010$
 Figure S.3: Example visualizations of annotated vs.
 predicted masks using four defect categories.
-Left: Input Image, Mid: Annotation, Right: Maskformer Predictions (SAM2).
+Left: Input Image, Mid: Annotation, Right: Maskformer Prediction.
 Segments are color-coded as follows: Foreground Banana, Background Banana, Old Bruise, Old Scar, New Bruise, New Scar.